### v2

@@ 4.2 Known kernels as special cases of $\kappa_{++}$
In the special case $\frac{1}{2}\bm{Q}$, and hence $\bm{D}$, equals $\bm{I}$, direct substitution reduces (27) to the multivariate Mehler kernel [45], [53, Thm. 1]
@@
Another special case of interest for $\kappa_{++}$ is the limit $\bm{D}\downarrow\bm{0}$ understood as $d_{i}\downarrow 0\>\forall i=1,\ldots,n$. Our next result (proof in Appendix A.7) is that in this limit, the kernel (27) recovers the heat kernel (4). This will be useful in the next subsection.
@@ 5 Discussions
(iii) From Proposition 4.5(iii), $\bm{M}\succ\bm{0}$ and hence the exponential term in (27) has negative definite argument. The same is true for the heat kernel (4). Thus the product of the exponential terms in (31) is bounded in $\left[0,1\right]$. So $|\kappa_{+}(t_{0},\bm{Vx},t,\bm{z})\>\widehat{\varphi}_{0}\left(\bm{V}^{\top}\bm{z}\right)|=\kappa_{+}(t_{0},\bm{Vx},t,\bm{z})\>\widehat{\varphi}_{0}\left(\bm{V}^{\top}\bm{z}\right)\leq\gamma(t,t_{0})\widehat{\varphi}_{0}\left(\bm{V}^{\top}\bm{z}\right)$, where $\gamma(t,t_{0}):=(4\pi(t-t_{0}))^{-p/2}\allowbreak(2\pi)^{-(n-p)/2}\left(\det(\bm{D}_{\left[i_{1}:i_{n-p}\right]})\right)^{1/4}/\sqrt{\det\left(\sinh\left(2(t-t_{0})\sqrt{\bm{D}_{\left[i_{1}:i_{n-p}\right]}}\right)\right)}$. Because $\widehat{\varphi}_{0}$ is Lebesgue measurable (given), applying the dominated convergence theorem separately for $|\bm{x}|\rightarrow\infty$ with finite $t$, and for $t\rightarrow\infty$, concludes the proof. $\blacksquare$
@@
In Fig. 3 and Fig. 4, we illustrate the solution $\widehat{\varphi}(t,\bm{x})$ in (32) for $n=2$ dimensions with $t_{0}=0$, initial conditions $\widehat{\varphi}_{0}(\cdot)=1$ and $\widehat{\varphi}_{0}(\cdot)=\mathcal{N}(\bm{0},\bm{I})$, respectively. In both figures, the columns correspond to different times and the rows correspond to different problem data: $\bm{Q}\succeq\bm{0}$. These results agree with the properties in Proposition 5.1, and were obtained by using the derived kernels in Theorem 4.3 and Theorem 4.7. For the two choices of $\widehat{\varphi}_{0}$ reported here, we analytically performed the integration in (32) using Lemma 2.3.
@@
$=\dfrac{\!\exp\!\left(\!-\frac{1}{2}\bm{x}^{\top}\sqrt{\bm{Q}}\tanh\left(2t\sqrt{\bm{Q}}\right)\bm{x}\!\right)}{\sqrt{\det\left(\cosh\left(2t\sqrt{\bm{Q}}\right)\right)}},$
@@
$\hskip 29.0pt\left.+\>{\rm{csch}}\left(2t\sqrt{\bm{D}}\right)\left(\sqrt{\bm{D}}\coth\left(2t\sqrt{\bm{D}}\right)+\bm{I}\right)^{-1}\sqrt{\bm{D}}\>{\rm{csch}}\left(2t\sqrt{\bm{D}}\right)\bigg{\}}\bm{y}\right),$
@@
$=\frac{(\det{\bm{Q}})^{\frac{1}{4}}}{\sqrt{\det\left(\sqrt{\bm{Q}}\cosh\left(2t\sqrt{\bm{Q}}\right)+\sinh\left(2t\sqrt{\bm{Q}}\right)\right)}}\!\exp\!\left(\!-\frac{1}{2}\bm{y}^{\top}\bm{Q}^{\frac{1}{2}}\bigg{\{}\coth\left(2t\sqrt{\bm{Q}}\right)\right.$
$\hskip 29.0pt\left.+\>{\rm{csch}}\left(2t\sqrt{\bm{Q}}\right)\left(\sqrt{\bm{Q}}\coth\left(2t\sqrt{\bm{Q}}\right)+\bm{I}\right)^{-1}\sqrt{\bm{Q}}\>{\rm{csch}}\left(2t\sqrt{\bm{Q}}\right)\bigg{\}}\bm{y}\right),$
@@
Future directions. We mention two directions in which our results could be extended. It will be of interest to extend our results for time-varying weights in quadratic state-cost-to-go, i.e., by considering matricial trajectory $\bm{Q}(t)$ that is continuous and bounded in $t$, and $\succeq\bm{0}$ for each $t\in[t_{0},t_{1}]$. Notice that such time-varying weights are straightforward to account for in classical LQ optimal control but it is not apparent how to generalize our Hermite polynomial-related developments for this case. From a control-theoretic viewpoint, it will also be of interest to solve problem (11) with the same quadratic state cost-to-go as in this work but with prior dynamics ${\rm{d}}\bm{x}_{t}=\bm{A}(t)\bm{x}_{t}{\rm{d}}t+\sqrt{2}\bm{B}(t){\rm{d}}\bm{w}_{t}$ as opposed to the case $\left(\bm{A}(t),\bm{B}(t)\right)=\left(\bm{0},\bm{I}\right)$ that we considered here following classical SB. These will be explored in our future work.
@@ Appendix A Proofs
Introducing a change of variable $y\mapsto\omega:=d^{1/4}y$ and letting $F(\omega):=f(d^{-1/4}\omega)$, we rewrite the above ODE as
$\sqrt{d}\frac{{\rm{d}}^{2}F}{{\rm{d}}\omega^{2}}-2d^{-1/4}\omega d^{1/4}\sqrt{d}\frac{{\rm{d}}F}{{\rm{d}}\omega}-\left(c+\sqrt{d}\right)F=0.$
@@
(49) $\frac{{\rm{d}}^{2}F}{{\rm{d}}\omega^{2}}-2\omega\frac{{\rm{d}}F}{{\rm{d}}\omega}-\left(\frac{c}{\sqrt{d}}+1\right)F=0,$
which is the Hermite’s ODE [20, p. 328-329].
Assuming $f$, and thus $F$, must be polynomially bounded, the solutions to (49) is $F(\omega)=H_{n}(\omega)$, the Hermite polynomials of order $n=-\frac{c}{2\sqrt{d}}-\frac{1}{2}\in\mathbb{N}_{0}$. Therefore, the solutions to (21) can be written as $Y(y)=a\exp{\left(-\frac{y^{2}\sqrt{d}}{2}\right)}H_{n}\left(d^{1/4}y\right)$ for some constant $a\in\mathbb{R}$ and $n=-\frac{c}{2\sqrt{d}}-\frac{1}{2}\in\mathbb{N}_{0}$. $\blacksquare$
@@
$H_{r}(cx)=\frac{1}{c^{r}}e^{(cx)^{2}}\left(-\frac{{\rm{d}}}{{\rm{d}}x}\right)^{r}\left(e^{-(cx)^{2}}\right)=(-2\iota)^{r}\frac{1}{\sqrt{\pi}}\int_{0}^{\infty}e^{-(\theta-cx\iota)^{2}}\theta^{r}{\rm{d}}\theta,\quad r\in\mathbb{N}_{0},$
@@
(54) $e^{-x^{2}}=\int e^{-\theta^{2}+2\iota x\theta}{\rm{d}}\theta/\sqrt{\pi}.$
@@
×$\qquad\qquad\qquad\qquad\times(-2\iota)^{m_{i}}\frac{1}{\sqrt{\pi}}\int_{t_{0}}^{\infty}e^{-\left(\theta-t_{0}-d_{i}^{1/4}z_{i}\iota\right)^{2}}\left(\theta-t_{0}\right)^{m_{i}}{\rm{d}}\theta$
$=\frac{e^{\left(d_{i}^{1/4}y_{i}\right)^{2}}}{\sqrt{\pi}}\int_{t_{0}}^{\infty}\sum_{m_{i}=0}^{\infty}\left(\frac{(\varrho_{i}\iota\left(\theta-t_{0}\right))^{m_{i}}}{\left(m_{i}!\right)\left(d_{i}^{1/4}\right)^{m_{i}}}\left(\frac{{\rm{d}}}{{\rm{d}}y_{i}}\right)^{m_{i}}\!\!e^{-\left(d_{i}^{1/4}y_{i}\right)^{2}}\!\right)e^{-\left(\theta-t_{0}-d_{i}^{1/4}z_{i}\iota\right)^{2}}{\rm{d}}\theta$
@@
$\sum_{m_{i}=0}^{\infty}\left(\frac{(\varrho_{i}\iota\left(\theta-t_{0}\right))^{m_{i}}}{\left(m_{i}!\right)\left(d_{i}^{1/4}\right)^{m_{i}}}\left(\frac{{\rm{d}}}{{\rm{d}}y_{i}}\right)^{m_{i}}\left(e^{-(d_{i}^{1/4}y_{i})^{2}}\right)\right)=e^{-\left(d_{i}^{1/4}y_{i}+\iota\varrho_{i}\left(\theta-t_{0}\right)\right)^{2}},$
@@
(55) $\dfrac{e^{d_{i}^{1/2}z_{i}^{2}}}{\sqrt{\pi}}\int_{t_{0}}^{\infty}\exp\left(-(1-\varrho_{i}^{2})(\theta-t_{0})^{2}+2d_{i}^{1/4}(\theta-t_{0})\iota\left(z_{i}-y_{i}\varrho_{i}\right)\right){\rm{d}}\theta.$
@@
()($\heartsuit$)=( ♡ ) = $\dfrac{e^{d_{i}^{1/2}z_{i}^{2}}}{\sqrt{1-\varrho_{i}^{2}}\sqrt{\pi}}\int_{0}^{\infty}\exp\left(-\sigma^{2}+2\iota\sigma\dfrac{d_{i}^{1/4}\left(z_{i}-y_{i}\varrho_{i}\right)}{\sqrt{1-\varrho_{i}^{2}}}\right){\rm{d}}\sigma$
@@
Letting $\bm{M}$ as in (28) and noting that $\det{\left(\bm{M}\right)}=\prod_{i=1}^{n}d_{i}>0$, (58) simplifies to (27). $\blacksquare$
@@
(ii) Thanks to the hyperbolic function identity $\coth^{2}(\cdot)-{\rm{csch}}^{2}(\cdot)=1$, definition (8) is satisfied, i.e., $\left(\bm{M}_{1}\bm{M}_{2}\right)^{\top}\bm{J}\allowbreak\left(\bm{M}_{1}\bm{M}_{2}\right)=\bm{J}$. Thus $\bm{M}_{1}\bm{M}_{2}\in{\rm{Sp}}\left(2n,\mathbb{R}\right)$.
To see positive definiteness, first note that $\bm{M}_{1}\succ 0$ for $\bm{D}\succ 0$ since
$\begin{pmatrix}\bm{y}&\bm{z}\end{pmatrix}\bm{M}_{1}\begin{pmatrix}\bm{y}\\ \bm{z}\end{pmatrix}$ $=\begin{pmatrix}\bm{y}&\bm{z}\end{pmatrix}\begin{bmatrix}\cosh\left(2t\sqrt{\bm{D}}\right)&-\bm{I}\\ -\bm{I}&\cosh\left(2t\sqrt{\bm{D}}\right)\end{bmatrix}\begin{pmatrix}\bm{y}\\ \bm{z}\end{pmatrix}$
$=\bm{y}^{\top}\underbrace{\cosh\left(2t\sqrt{\bm{D}}\right)}_{\succ\bm{I}}\bm{y}-2\langle\bm{y},\bm{z}\rangle+\bm{z}^{\top}\underbrace{\cosh\left(2t\sqrt{\bm{D}}\right)}_{\succ\bm{I}}\bm{z}$
@@
Also, $\bm{D}\succ\bm{0}$ implies the positive diagonal matrix $\bm{M}_{2}\succ\bm{0}$, and therefore,
${\rm{eig}}\left(\bm{M}_{1}\bm{M}_{2}\right)={\rm{eig}}\left(\bm{M}_{1}\bm{M}_{2}^{1/2}\bm{M}_{2}^{1/2}\right)={\rm{eig}}\left(\underbrace{\bm{M}_{2}^{1/2}\bm{M}_{1}\bm{M}_{2}^{1/2}}_{\text{congruence transform of $\bm{M}_{1}$}}\right)>0.$
Additionally, $\bm{M}_{1}\bm{M}_{2}$ being symmetric, we have that $\bm{M}_{1}\bm{M}_{2}\succ\bm{0}$.
@@
${\rm{eig}}\left(\bm{M}\right)={\rm{eig}}\left(\underbrace{\begin{bmatrix}\bm{D}^{1/4}&\bm{0}\\ \bm{0}&\bm{D}^{1/4}\end{bmatrix}\bm{M}_{1}\bm{M}_{2}\begin{bmatrix}\bm{D}^{1/4}&\bm{0}\\ \bm{0}&\bm{D}^{1/4}\end{bmatrix}}_{\text{congruence transform of $\bm{M}_{1}\bm{M}_{2}$}}\right)>0,$
and $\bm{M}$ is symmetric, together implies $\bm{M}\succ\bm{0}$. $\blacksquare$
A.7 Proof for Theorem 4.6

### v4

@@ 4.2 Known kernels as special cases of $\kappa_{++}$
In the special case $\frac{1}{2}\bm{Q}$, and hence $\bm{D}$, equals $\bm{I}$, direct substitution reduces (27) to the multivariate Mehler kernel [48], [56, Thm. 1]
@@
Another special case of interest for $\kappa_{++}$ is the limit $\bm{D}\downarrow\bm{0}$ understood as $d_{i}\downarrow 0\>\forall i=1,\ldots,n$. Our next result (proof in Appendix A.8) is that in this limit, the kernel (27) recovers the heat kernel (4). This will be useful in subsection 4.3.
@@
That $\kappa_{0}$ arises as a limiting case of $\kappa_{++}$ is consistent with the principle of least action in Proposition 4.6 since in the limit $\bm{D}\downarrow\bm{0}$, the Lagrangian in (29) equals $\frac{1}{2}|\bm{v}|^{2}$. Then, (30) yields the squared Euclidean distance $|\bm{y}-\bm{z}|^{2}$ which is what appears inside the exponential in (4) up to the usual scaling.
@@ 5 Discussions
(iii) From Proposition 4.5(iii), $\bm{M}\succ\bm{0}$ and hence the exponential term in (27) has negative definite argument. The same is true for the heat kernel (4). Thus the product of the exponential terms in (33) is bounded in $\left[0,1\right]$. So $|\kappa_{+}(t_{0},\bm{Vx},t,\bm{z})\>\widehat{\varphi}_{0}\left(\bm{V}^{\top}\bm{z}\right)|=\kappa_{+}(t_{0},\bm{Vx},t,\bm{z})\>\widehat{\varphi}_{0}\left(\bm{V}^{\top}\bm{z}\right)\leq\zeta(t,t_{0})\widehat{\varphi}_{0}\left(\bm{V}^{\top}\bm{z}\right)$, where $\zeta(t,t_{0}):=(4\pi(t-t_{0}))^{-p/2}\allowbreak(2\pi)^{-(n-p)/2}\left(\det(\bm{D}_{\left[i_{1}:i_{n-p}\right]})\right)^{1/4}/\sqrt{\det\left(\sinh\left(2(t-t_{0})\sqrt{\bm{D}_{\left[i_{1}:i_{n-p}\right]}}\right)\right)}$. Because $\widehat{\varphi}_{0}$ is Lebesgue measurable (given), applying the dominated convergence theorem separately for $|\bm{x}|\rightarrow\infty$ with finite $t$, and for $t\rightarrow\infty$, concludes the proof. $\blacksquare$
@@
Figure 5: Solution to PDE (12a) in $n=2$ dimensions with $t_{0}=0$ and initial condition $\widehat{\varphi}_{0}(\cdot)=\rho_{\infty}$. All subfigures are over the spatial domain $\left[-0.25,1.25\right]^{2}$.
In Fig. 3, Fig. 4 and Fig. 5, we illustrate the solution $\widehat{\varphi}(t,\bm{x})$ in (34) for $n=2$ dimensions with $t_{0}=0$, initial conditions $\widehat{\varphi}_{0}(\cdot)=1$, $\widehat{\varphi}_{0}(\cdot)=\mathcal{N}(\bm{0},\bm{I})$, and $\widehat{\varphi}_{0}(\cdot)\propto\exp\left(-f(x_{1},x_{2})/35\right)$ where $f(x_{1},x_{2}):=\left((10x_{1}-5)^{2}+10x_{2}-16\right)^{2}+\left(10x_{1}-12+(10x_{2}-5)^{2}\right)^{2}$, $(x_{1},x_{2})\in[0,1]^{2}$, respectively. We point out that the last $\widehat{\varphi}_{0}$ is of non-Gaussian Gibbs-type [34, Sec. V-C], and $f$ there is a scaled Himmelblau function, often used as a benchmark in nonconvex optimization [36].
In these figures, the columns correspond to different times and the rows correspond to different problem data: $\bm{Q}\succeq\bm{0}$. These results agree with the properties in Proposition 5.1, and were obtained by using the derived kernels in Theorem 4.3 and Theorem 4.9. For the three cases of $\widehat{\varphi}_{0}$ reported here, we analytically performed the integration in (34) using Lemma 2.3.
@@
$=\dfrac{\!\exp\!\left(\!-{\frac{1}{4}}\bm{x}^{\top}\sqrt{2\bm{Q}}\tanh\left(t\sqrt{2\bm{Q}}\right)\bm{x}\!\right)}{\sqrt{\det\left(\cosh\left(t\sqrt{2\bm{Q}}\right)\right)}},$
@@
$\hskip 29.0pt\left.-\>{\rm{csch}}\left(2t\sqrt{\bm{D}}\right)\left(\sqrt{\bm{D}}\coth\left(2t\sqrt{\bm{D}}\right)+\bm{I}\right)^{-1}\sqrt{\bm{D}}\>{\rm{csch}}\left(2t\sqrt{\bm{D}}\right)\bigg{\}}\bm{y}\right),$
@@
$={\color[rgb]{0,0,0}{\frac{2^{-n/4}(\det\bm{Q})^{\frac{1}{4}}}{\sqrt{\det\left(\sqrt{2\bm{Q}}\cosh\left(t\sqrt{2\bm{Q}}\right)/2+\sinh\left(t\sqrt{2\bm{Q}}\right)\right)}}\!\exp\!\left(\!-\frac{1}{4}\bm{x}^{\top}\sqrt{2\bm{Q}}\bigg{\{}\coth\left(t\sqrt{2\bm{Q}}\right)\right.}}$
$\hskip 29.0pt{\color[rgb]{0,0,0}{\left.-\>{\rm{csch}}\left(t\sqrt{2\bm{Q}}\right)\left(\frac{1}{4}\sqrt{2\bm{Q}}\coth\left(t\sqrt{2\bm{Q}}\right)+\bm{I}\right)^{-1}\sqrt{2\bm{Q}}\>{\rm{csch}}\left(t\sqrt{2\bm{Q}}\right)\bigg{\}}\bm{x}\right)}},$
@@
Likewise, the results in Fig. 5 show the transient action of the kernel on non-Gaussian $\widehat{\varphi}_{0}$.
@@
Future directions. We mention several directions in which our results could be extended. It will be of interest to extend our results for time-varying weights in quadratic state-cost-to-go, i.e., by considering matricial trajectory $\bm{Q}(t)$ that is continuous and bounded in $t$, and $\succeq\bm{0}$ for each $t\in[t_{0},t_{1}]$. Such time-varying weights are straightforward to account for in classical LQ optimal control but it is not apparent how to generalize our Hermite polynomial-related developments for this case. From a control-theoretic viewpoint, it will also be of interest to solve problem (11) with the same quadratic state cost-to-go as in this work but with prior dynamics ${\rm{d}}\bm{x}_{t}=\bm{A}(t)\bm{x}_{t}{\rm{d}}t+\sqrt{2}\bm{B}(t){\rm{d}}\bm{w}_{t}$ as opposed to the case $\left(\bm{A}(t),\bm{B}(t)\right)=\left(\bm{0},\bm{I}\right)$ that we considered here following classical SB. In a different vein, it will be of interest to find the worst-case contraction coefficient [64] for the recursion in Fig. 2 as function of the matrix $\bm{Q}$, using the derived kernel. These will be explored in our future work.
@@ Appendix A Proofs
Introducing a change of variable $y\mapsto\xi:=d^{1/4}y$ and letting $F(\xi):=f(d^{-1/4}\xi)$, we rewrite the above ODE as
$\sqrt{d}\frac{{\rm{d}}^{2}F}{{\rm{d}}\xi^{2}}-2d^{-1/4}\xi d^{1/4}\sqrt{d}\frac{{\rm{d}}F}{{\rm{d}}\xi}-\left(c+\sqrt{d}\right)F=0.$
@@
(51) $\frac{{\rm{d}}^{2}F}{{\rm{d}}\xi^{2}}-2\xi\frac{{\rm{d}}F}{{\rm{d}}\xi}-\left(\frac{c}{\sqrt{d}}+1\right)F=0,$
which is the Hermite’s ODE [21, p. 328-329].
Assuming $f$, and thus $F$, must be polynomially bounded, the solutions to (51) is $F(\xi)=H_{n}(\xi)$, the Hermite polynomials of order $n=-\frac{c}{2\sqrt{d}}-\frac{1}{2}\in\mathbb{N}_{0}$. Therefore, the solutions to (21) can be written as $Y(y)=a\exp{\left(-\frac{y^{2}\sqrt{d}}{2}\right)}H_{n}\left(d^{1/4}y\right)$ for some constant $a\in\mathbb{R}$ and $n=-\frac{c}{2\sqrt{d}}-\frac{1}{2}\in\mathbb{N}_{0}$. $\blacksquare$
@@
$H_{r}(cx)=\frac{1}{c^{r}}e^{(cx)^{2}}\left(-\frac{{\rm{d}}}{{\rm{d}}x}\right)^{r}\left(e^{-(cx)^{2}}\right)=(-2{\mathrm{i}})^{r}\frac{1}{\sqrt{\pi}}\int_{0}^{\infty}e^{-(\theta-cx{\mathrm{i}})^{2}}\theta^{r}{\rm{d}}\theta,\quad r\in\mathbb{N}_{0},$
@@
(56) $e^{-x^{2}}=\int e^{-\theta^{2}+2{\mathrm{i}}x\theta}{\rm{d}}\theta/\sqrt{\pi}.$
@@
×$\qquad\qquad\qquad\qquad\times(-2{\mathrm{i}})^{m_{i}}\frac{1}{\sqrt{\pi}}\int_{t_{0}}^{\infty}e^{-\left(\theta-t_{0}-d_{i}^{1/4}z_{i}{\mathrm{i}}\right)^{2}}\left(\theta-t_{0}\right)^{m_{i}}{\rm{d}}\theta$
$=\frac{e^{\left(d_{i}^{1/4}y_{i}\right)^{2}}}{\sqrt{\pi}}\int_{t_{0}}^{\infty}\sum_{m_{i}=0}^{\infty}\left(\frac{(\varrho_{i}{\mathrm{i}}\left(\theta-t_{0}\right))^{m_{i}}}{\left(m_{i}!\right)\left(d_{i}^{1/4}\right)^{m_{i}}}\left(\frac{{\rm{d}}}{{\rm{d}}y_{i}}\right)^{m_{i}}\!\!e^{-\left(d_{i}^{1/4}y_{i}\right)^{2}}\!\right)e^{-\left(\theta-t_{0}-d_{i}^{1/4}z_{i}{\mathrm{i}}\right)^{2}}{\rm{d}}\theta$
@@
$\sum_{m_{i}=0}^{\infty}\left(\frac{(\varrho_{i}{\mathrm{i}}\left(\theta-t_{0}\right))^{m_{i}}}{\left(m_{i}!\right)\left(d_{i}^{1/4}\right)^{m_{i}}}\left(\frac{{\rm{d}}}{{\rm{d}}y_{i}}\right)^{m_{i}}\left(e^{-(d_{i}^{1/4}y_{i})^{2}}\right)\right)=e^{-\left(d_{i}^{1/4}y_{i}+{\mathrm{i}}\varrho_{i}\left(\theta-t_{0}\right)\right)^{2}},$
@@
(57) $\dfrac{e^{d_{i}^{1/2}z_{i}^{2}}}{\sqrt{\pi}}\int_{t_{0}}^{\infty}\exp\left(-(1-\varrho_{i}^{2})(\theta-t_{0})^{2}+2d_{i}^{1/4}(\theta-t_{0}){\mathrm{i}}\left(z_{i}-y_{i}\varrho_{i}\right)\right){\rm{d}}\theta.$
@@
()($\heartsuit$)=( ♡ ) = $\dfrac{e^{d_{i}^{1/2}z_{i}^{2}}}{\sqrt{1-\varrho_{i}^{2}}\sqrt{\pi}}\int_{0}^{\infty}\exp\left(-\sigma^{2}+2{\mathrm{i}}\sigma\dfrac{d_{i}^{1/4}\left(z_{i}-y_{i}\varrho_{i}\right)}{\sqrt{1-\varrho_{i}^{2}}}\right){\rm{d}}\sigma$
@@
Letting ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}$ as in (28) and noting that $\det{\left({\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}\right)}=\prod_{i=1}^{n}d_{i}>0$, (60) simplifies to (27). $\blacksquare$
@@
(ii) Thanks to the hyperbolic function identity $\coth^{2}(\cdot)-{\rm{csch}}^{2}(\cdot)=1$, definition (8) is satisfied, i.e., $\left({\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}\bm{M}_{tt_{0}}^{(2)}}}\right)^{\top}\bm{J}\allowbreak\left({\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}\bm{M}_{tt_{0}}^{(2)}}}\right)=\bm{J}$. Thus ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}\bm{M}_{tt_{0}}^{(2)}}}\in{\rm{Sp}}\left(2n,\mathbb{R}\right)$.
To see positive definiteness, first note that ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}}}\succ 0$ for $\bm{D}\succ 0$ since
$\begin{pmatrix}\bm{y}&\bm{z}\end{pmatrix}{\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}}}\begin{pmatrix}\bm{y}\\ \bm{z}\end{pmatrix}$ $=\begin{pmatrix}\bm{y}&\bm{z}\end{pmatrix}\begin{bmatrix}\cosh\left(2{\color[rgb]{0,0,0}{\left(t-t_{0}\right)}}\sqrt{\bm{D}}\right)&-\bm{I}\\ -\bm{I}&\cosh\left(2{\color[rgb]{0,0,0}{\left(t-t_{0}\right)}}\sqrt{\bm{D}}\right)\end{bmatrix}\begin{pmatrix}\bm{y}\\ \bm{z}\end{pmatrix}$
$=\bm{y}^{\top}\underbrace{\cosh\left(2{\color[rgb]{0,0,0}{\left(t-t_{0}\right)}}\sqrt{\bm{D}}\right)}_{\succ\bm{I}}\bm{y}-2\langle\bm{y},\bm{z}\rangle+\bm{z}^{\top}\underbrace{\cosh\left(2{\color[rgb]{0,0,0}{\left(t-t_{0}\right)}}\sqrt{\bm{D}}\right)}_{\succ\bm{I}}\bm{z}$
@@
Also, $\bm{D}\succ\bm{0}$ implies the positive diagonal matrix ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(2)}}}\succ\bm{0}$, and therefore,
${\rm{eig}}\left({\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}\bm{M}_{tt_{0}}^{(2)}}}\right)$ $={\rm{eig}}\left({\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}\left(\bm{M}_{tt_{0}}^{(2)}\right)^{1/2}\left(\bm{M}_{tt_{0}}^{(2)}\right)^{1/2}}}\right)$
$={\rm{eig}}\left(\underbrace{{\color[rgb]{0,0,0}{\left(\bm{M}_{tt_{0}}^{(2)}\right)^{1/2}\bm{M}_{tt_{0}}^{(1)}\left(\bm{M}_{tt_{0}}^{(2)}\right)^{1/2}}}}_{\text{congruence transform of ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}}}$}}\right)>0.$
Additionally, ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}\bm{M}_{tt_{0}}^{(2)}}}$ being symmetric, we have that ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}\bm{M}_{tt_{0}}^{(2)}}}\succ\bm{0}$.
@@
${\rm{eig}}\left({\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}\right)={\rm{eig}}\left(\underbrace{\begin{bmatrix}\bm{D}^{1/4}&\bm{0}\\ \bm{0}&\bm{D}^{1/4}\end{bmatrix}{\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}\bm{M}_{tt_{0}}^{(2)}}}\begin{bmatrix}\bm{D}^{1/4}&\bm{0}\\ \bm{0}&\bm{D}^{1/4}\end{bmatrix}}_{\text{congruence transform of ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}^{(1)}\bm{M}_{tt_{0}}^{(2)}}}$}}\right)>0,$
and ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}$ is symmetric, together implies ${\color[rgb]{0,0,0}{\bm{M}_{tt_{0}}}}\succ\bm{0}$. $\blacksquare$
A.7 Proof for Proposition 4.6
That the minimum in (30) is attained for the $L$ in (29), follows from the strict convexity and coercivity of $\bm{v}\mapsto L\left(\cdot,\bm{v}\right)$.
For notational convenience, let $\omega_{i}:=2\sqrt{d_{i}}$ $\forall i=1,\ldots,n$. The optimal solution $\bm{q}_{\rm{opt}}(\tau)$, $\tau\in[t_{0},t]$, solves the Euler-Lagrange equation $\frac{{\rm{d}}^{2}}{{\rm{d}}\tau^{2}}{\bm{q}}^{\rm{opt}}=4\bm{Dq}^{\rm{opt}}$, or in component form: $\frac{{\rm{d}}^{2}}{{\rm{d}}\tau^{2}}{q_{i}}^{\rm{opt}}=\omega_{i}^{2}q_{i}^{\rm{opt}}$ subject to the endpoint constraints $q_{i}^{\rm{opt}}(t_{0})=y_{i}$, $q_{i}^{\rm{opt}}(t)=z_{i}$. The general solution is $q_{i}^{\rm{opt}}(\tau)=a_{i}e^{\omega_{i}\tau}+b_{i}e^{-\omega_{i}\tau}$, and the endpoint constraints determine the constants $\left(a_{i},b_{i}\right)$ as
(61) $\left(a_{i},b_{i}\right)=\left(-y_{i}e^{-\omega_{i}t}+z_{i}e^{-\omega_{i}t_{0}},\>y_{i}e^{\omega_{i}t}-z_{i}e^{\omega_{i}t_{0}}\right)/2\sinh\left(\omega_{i}\left(t-t_{0}\right)\right).$
Then, (30) gives
${\rm{dist}}_{tt_{0}}^{2}(\bm{y},\bm{z})$ $=\int_{t_{0}}^{t}\sum_{i=1}^{n}\left(\dfrac{1}{2}\left(\dfrac{{\rm{d}}}{{\rm{d}}\tau}q_{i}^{\rm{opt}}(\tau)\right)^{2}+\dfrac{1}{2}\omega_{i}^{2}\left(q_{i}^{\rm{opt}}(\tau)\right)^{2}\right){\rm{d}}\tau$
$=\displaystyle\sum_{i=1}^{n}\omega_{i}^{2}\int_{t_{0}}^{t}\left(a_{i}^{2}e^{2\omega_{i}\tau}+b_{i}^{2}e^{-2\omega_{i}\tau}\right){\rm{d}}\tau$
$=\sum_{i=1}^{n}\dfrac{\omega_{i}}{2}\big{\{}a_{i}^{2}\left(e^{2\omega_{i}t}-e^{2\omega_{i}t_{0}}\right)-b_{i}^{2}\left(e^{-2\omega_{i}t}-e^{-2\omega_{i}t_{0}}\right)\big{\}}$
(62) $=\sum_{i=1}^{n}\dfrac{\omega_{i}}{2}\dfrac{\left(y_{i}^{2}+z_{i}^{2}\right)\cosh\left(\omega_{i}\left(t-t_{0}\right)\right)-2y_{i}z_{i}}{\sinh\left(\omega_{i}\left(t-t_{0}\right)\right)},$
where in the last line, we substituted for $(a_{i},b_{i})$ from (61), and simplified the resulting expression using the identity $\sinh\left(2\left(\cdot\right)\right)=2\sinh\left(\cdot\right)\cosh\left(\cdot\right)$.
Since $\omega_{i}=2\sqrt{d_{i}}$ $\forall i=1,\ldots,n$, we see that (62) is indeed the desired expression, c.f. the expression inside the exponential in (60). This completes the proof. $\blacksquare$
A.8 Proof for Theorem 4.8